OpenStudy (anonymous):

A cubic polynomial F has three "different" real zeros a,b and c. The coefficient of x^3 is positive. Show that F`(a)+F`(b)+F`(c) > 0

OpenStudy (anonymous):

\[F(x)=p(x-a)(x-b)(x-c)\] is probably a good start, where \(p>0\) take the derivative via the product rule and see what you get i didn't actually do it, but is seems like a start to me

OpenStudy (anonymous):

Thanks, would that be all I need to do?

OpenStudy (anonymous):

i don't know i am writing it now but it seems like the only way to start

OpenStudy (anonymous):

\[F'(x)=p(x-b)(x-c)+p(x-a)(x-b)+p(x-a)(x-c)\] is a start then compute \(F'(a)\) etc. add. and see what you get

OpenStudy (anonymous):

Not sure how to do that either, I've only just begun attempting this type of problems.

OpenStudy (anonymous):

All help is very appreciated

OpenStudy (anonymous):

man this is not nearly as easy as i though

OpenStudy (anonymous):

lets ignore the \(p\) for a second, as it is positive, lets say it is 1

OpenStudy (anonymous):

\[F'(x)=(x-b)(x-c)+(x-a)(x-b)+(x-a)(x-c)\] \[F'(a)=(a-b)(a-c)\] \[F'(b)=(b-a)(b-c)\] \[F'(c)=(c-a)(c-b)\]

OpenStudy (anonymous):

add them up and get \[(a-b)(a-c)+(b-a)(b-c)+(c-a)(c-b)\]

OpenStudy (anonymous):

there might be some elementary reason why this is positive, but i don't see it if you do a raft of algebra you can write this as the sum of two squares, and so it has to be positive (non negative actually)

OpenStudy (anonymous):

oh actually positive, since it is only zero if \(a=b\) or \(b=c\) and you are told that they are different

OpenStudy (anonymous):

@Zarkon do you see right of the bat that this must be positive?

OpenStudy (zarkon):

look at \[\frac{1}{2}[(a-b)^2+(a-c)^2+(b-c)^2]\]

OpenStudy (anonymous):

*off

OpenStudy (zarkon):

\[f'(a)+f'(b)+f'(c)=\frac{1}{2}[(a-b)^2+(a-c)^2+(b-c)^2]\]

OpenStudy (zarkon):

my f being your F ;)

OpenStudy (zarkon):

suppressing the p

OpenStudy (anonymous):

i believe you but i have to figure out why

OpenStudy (anonymous):

Same here

OpenStudy (anonymous):

oh, algebra is why

OpenStudy (anonymous):

Heh, could you explain it? I feel as though I am not smart enough to figure it out by myself at the moment :/

OpenStudy (anonymous):

it is algebra: multiply out, combine like terms, and see what you get although how @Zarkon recognized it as \[\frac{1}{2}\left((a-b)^2+(a-c)^2+(b-c)^2\right)\] is for him to know

OpenStudy (anonymous):

Does that prove that F`(a)+F`(b)+F`(c) > 0 is possible?

OpenStudy (anonymous):

a square is always positive unless it is zero

OpenStudy (zarkon):

it was a guess

OpenStudy (anonymous):

really? damn! good guess

OpenStudy (zarkon):

I thought it might be that ...so I expanded both

OpenStudy (anonymous):

@Dahlioz multiply it out and see that it works

OpenStudy (anonymous):

i cheated and used wolfram, probably would have been better to do it by hand, which i did afterwards

OpenStudy (anonymous):

wolf gives this http://www.wolframalpha.com/input/?i=%28a-b%29%28a-c%29%2B%28b-a%29%28b-c%29%2B%28c-a%29%28c-b%29 although not \[\frac{1}{4}\left(2a-b-c\right)^2+\frac{3}{4}\left(b-c\right)^2\] which also gives it

OpenStudy (zarkon):

ah...cool

OpenStudy (anonymous):

not nearly as cool as the sum of three squares

OpenStudy (zarkon):

I agree ;)

OpenStudy (anonymous):

Thanks guys, you both have been very helpful

OpenStudy (anonymous):

Just one question, about what grade/year of school would this be suitable for?

OpenStudy (anonymous):

But why can't (2a-b-c) be equal to 0?

OpenStudy (anonymous):

As wolfram gives 1/4 (2 a-b-c)^2+3/4 (b^2-2 b c+c^2) rather than 1/4(2a−b−c)^2+3/4(b−c)^2. @Zarkon Sorry for disturbing you guys with this.

OpenStudy (zarkon):

those two expressions are equal

OpenStudy (anonymous):

Ahh, I see